Delta Enterprises, Inc. has a WACC of 10% and is considering a project that requires a cash outlay of $1,250 now with cash inflows of $500 at the end of each year for the next 5 years. What is the project’s Discounted Payback? Enter your answer rounded to two decimal places.
Answer #1

Present value of year 1 cash flow = 500 / (1 + 0.1) = 454.545455

Present value of year 2 cash flow = 500 / (1 + 0.1)2 = 413.22314

Present value of year 3 cash flow = 500 / (1 + 0.1)3 = 375.6574

Present value of year 4 cash flow = 500 / (1 + 0.1)4 = 341.506728

Present value of year 5 cash flow = 500 / (1 + 0.1)5 = 310.460662

Cumulative cash flow for year 0 = -1,250

Cumulative cash flow for year 1 = -1,250 + 454.545455 = -795.454545

Cumulative cash flow for year 2 = -795.454545 + 413.22314 = -382.231405

Cumulative cash flow for year 3 = -382.231405 + 375.6574 = -6.574005

Cumulative cash flow for year 4 = -6.574005 + 341.506728 = 334.93

6.574005 / 341.506728 = 0.02

Discounted payback = 3 + 0.02 = 3.02 years